# The One RING [Uncaptioned image]: a Robotic Indoor Navigation Generalist

### v1

@@ Abstract
Modern robots vary significantly in shape, size, and sensor configurations used to perceive and interact with their environments. However, most navigation policies are embodiment-specific; a policy learned using one robot’s configuration does not typically gracefully generalize to another. Even small changes in the body size or camera viewpoint may cause failures. With the recent surge in custom hardware developments, it is necessary to learn a single policy that can be transferred to other embodiments, eliminating the need to (re-)train for each specific robot. In this paper, we introduce Ring (Robotic Indoor Navigation Generalist), an embodiment-agnostic policy, trained solely in simulation with diverse randomly initialized embodiments at scale. Specifically, we augment the AI2-THOR simulator with the ability to instantiate robot embodiments with controllable configurations, varying across body size, rotation pivot point, and camera configurations. In the visual object-goal navigation task, Ring achieves robust performance on real unseen robot platforms (Stretch RE-1, LoCoBot, Unitree’s Go1), achieving an average of $72.1\%$ and $78.9\%$ success rate across $5$ embodiments in simulation and $4$ robot platforms in the real world.
Figure 1: (A) We train on one million randomly generated embodiments in simulation varying camera configurations, body size, and rotation pivot point. The plot shows the t-SNE visualization of embodiment parameters $\mathbf{c}_{e}\in\mathbb{R}^{19}$ for 30k random agents and three specific robots (robots are for visualization-we do not train on any real robot embodiment parameters). Egocentric views from the first camera are shown for 10 sample agents. (B) Our trained policy transfers zero-shot to a wide range of embodiments in the real-world including Stretch RE-1, LoCoBot, and Unitree Go1, as well as a human embodiment. (C) The Ring policy displays embodiment-adaptive behavior, adjusting its navigation strategy based on its embodiment.
@@ 1 Introduction
Robot embodiments are diverse and are constantly evolving to better suit new environments and tasks. This range in body configurations—differences in size, shape, wheeled or legged locomotion, and sensor configurations—not only shapes how robots perceive the world but also how they act in it. A robot with a wide field of view (FoV) or multiple cameras can scan its surroundings quickly, while one with a narrower view might need to more actively explore a room. A small robot can squeeze through tight spaces, a low-profile one can duck under furniture, and a larger robot may need to follow more conservative routes. The influence of embodiment on behavior means a policy trained on one design, or even several, often does not perform well out of domain [62].
@@
There has been progress towards scalable cross-embodiment training [34, 45] and in developing general-purpose navigation policies [42, 43, 13, 56]. While these methods demonstrate some transfer to unseen embodiments, they require construction of topological maps or graphs and suffer performance degradation with relatively small changes in embodiment (e.g., camera position modification on the same robot). Potentially, this is due to these methods relying on the small amount of real-world data available in public datasets-only around 20 embodiments in total. This highlights the need for a more comprehensive solution that reliably covers the wide range of possible embodiments without retraining or additional adaptation.
@@
We introduce Ring, a Robotic Indoor Navigation Generalist. Ring is trained exclusively in simulation, without any use of real-world robot embodiments. In other words, all robot platforms we evaluate on (i.e., Stretch RE-1, LoCoBot, Unitree’s A1) are unseen by Ring during training. We leverage simulation to randomly sample 1 Million agent body configurations, varying the robot’s camera parameters, collider sizes, and center of rotation. Concretely, each embodiment consists of a collider box of varying dimensions and cameras with randomized FoV and dimensions, placed randomly within the collider box. Fig. 1-A shows a t-SNE [46] visualization of body parameters for $30$k random agents in our generated data.
Our method is inspired by the recent success in real-world experiments while training at large-scale only in simulation [17, 62, 24]. Simulation training is able to benefit from the vast scale of scenes (150k ProcTHOR houses [12]) and objects (40k+ annotated 3D objects from Objaverse [11]) in the AI2-THOR simulator. Extensive domain randomization on visual observations and the use of pre-trained visual encoders then allows simulation-trained policies to bridge the sim-to-real gap. We follow the training procedure outlined in FLaRe [24], first training our policy on expert trajectories collected from $1$M randomized embodiments and subsequently fine-tuning it with on-policy reinforcement learning (RL) within the simulator.
@@
Our results demonstrate generalization to truly unseen embodiments. Ring generalizes to diverse real-world embodiments without any adaptation, despite being trained exclusively in simulation without access to the real robot configurations. We evaluate our policy in a zero-shot setting across a variety of embodiments, including the Stretch RE-1, LoCoBot, Unitree’s A1, and even “Navigation Assistants”, wherein a human user captures ego-centric observations on their phone and prompt Ring policy to predict actions to navigate. Ring achieves $72.1\%$ and $78.9\%$ success rate on average, outperforming the best baseline significantly, both in simulation and the real-world.
We highlight three key characteristics of Ring: 1) it displays zero-shot generalization to unseen embodiments, keeping a consistently high performance (Sec. 4.1); 2) it can zero-shot transfer to the real-world without any adaptation or real-world-specific finetuning. (Sec. 4.2); 3) it can be adapted to an embodiment-specialized policy with even better performance with minimal finetuning (Sec. 4.3); and 4) at inference, it dynamically adjusts its behavior based on the embodiment (Sec. 4.4, Fig. 6). Ring can be directly deployed to navigate any robot platform, is easy to install, and is ready for use by researchers in the community. We will release our pretrained models, generated data, and training code.
@@ 2 Related work
Cross-embodiment training has received substantial attention from the research community. Arguably the most representative of a large body of recent work [56, 13, 33, 47, 14, 30, 53, 19, 23, 7, 21, 3, 25, 28, 52, 60], Open-X-Embodiment (OXE) [9] is the fruit of a large collaboration to cover many robotic tasks, with special emphasis in manipulation. Its usage in RT-X results in a notable performance gain in emergent skill evaluations in comparison to RT-2 [5]. Despite the 1.5 million trajectories across 22 embodiments present in their dataset, the enormous cost of data collection in the real world makes further scaling challenging. CrossFormer [13] trains a transformer-based policy on 900k trajectories across 30 robots, including a subset of OXE, navigation data from GNM [42], manipulation data from DROID [27], and additional collected data. Due to the relatively sparse amount of embodiments observed during training and the target low-level control, it does not generalize to unseen embodiments. GET-zero [36] focuses on dexterous manipulation, and proposes to inform the policy with the structure of the embodiment via a connectivity graph to bias the attention. In contrast, we generate an arbitrarily large amount of embodiments for training our policy, enabling zero-shot deployment to new embodiments without accessing the embodiment structure.
Foundational navigation policies.
@@
Following the success in recent developments for point-goal navigation [50], locomotion [39, 4, 41], agile control [51], exploration [6, 55, 57], and social navigation [37], comparable results in more nuanced tasks like semantic or object-goal navigation (ObjectNav) [2, 15, 61, 31, 48, 40, 58, 26] remain elusive due to a lack of efficient exploration and semantic understanding capabilities. Recently, with powerful pretrained foundational vision models [35, 63] and large-scale procedurally generated virtual environments [11], notable progress in end-to-end ObjectNav policy learning for specific embodiments has been achieved by means of imitation learning (IL) from shortest-path trajectories [17], RL [62], or combinations thereof [24]. In image-goal navigation, NoMaD [44], which extends ViNT [43], uses a diffusion policy to control a single embodiment. With the same goal in mind, GNM [42] trains navigation policies across 6 embodiments using IL. In contrast, our policy benefits from finetuning with RL, improving resilience to compounding errors. Additionally, thanks to training with large-scale randomized embodiments in simulation, Ring learns a single policy to navigate any embodiment, generalizing to truly unseen robot platforms in the real world. Furthermore, NoMaD, ViNT, GNM, and Mobility VLA [54] all require topological map or graph reconstruction for high-level planning, whereas our policy is fully end-to-end and explores novel scenes without an explicit map. While several efforts [32, 59, 1] focus on learning embodiment-agnostic policies using LLMs or VLMs, they address only short-horizon navigation tasks and perform single-step predictions. In contrast, Ring models temporal information through a transformer decoder.
@@ 3 Ring
Table 1: Random Embodiment Parameters. We generate 1M different embodiments sampled from the ranges above.
With the growing diversity of robots used in research labs and real-world applications, there remains a need for a policy that can operate a wide range of embodiments and transfer, in a zero- or few-shot manner, to unseen robots. We introduce Ring, a generalist policy for indoor visual navigation that learns from a broad spectrum of embodiments, trained exclusively in simulation, without any direct use of actual robot embodiments. We show that training on an extensive range of ${\sim}1$M random embodiments results in a robust navigation policy, enabling zero-shot transfer to unseen real-world robot embodiments. To train Ring, we define the space of random embodiments (Sec. 3.2), enable generation of expert trajectories for random embodiments in simulation (see Appendix 7), and use state-of-the-art architecture designs (Sec. 3.3) to train with a combination of IL and RL methods (Sec. 3.4).
3.1 Problem formulation
Learning a navigation policy across multiple embodiments is a multi-task robotic problem. We define the space of possible embodiments as $\mathcal{E}$, where each embodiment $e\in\mathcal{E}$ is characterized by a configuration vector $\mathbf{c}_{e}$, which includes parameters such as camera settings, agent collider size, center of rotation, etc. Each task can be modeled as a Partially Observable Markov Decision Process (POMDP), denoted as $(\mathcal{S},\mathcal{A},\mathcal{E},\mathcal{O}_{e},\mathcal{T}_{e},R,\mathcal{L},P(s_{0}),\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ are the state and action spaces. The observation space $\mathcal{O}_{e}$ varies across embodiments due to differences in camera parameters and sensor configurations. The observation at time $t$ for embodiment $e$, $o_{t}^{e}=\mathcal{O}_{e}(s_{t},\mathbf{c}_{e})$, is a function of both the state $s_{t}$ and embodiment parameters $\mathbf{c}_{e}$. Given an action $a_{t}$, the next state follows the transition dynamics $s_{t+1}\sim\mathcal{T}_{e}(s_{t+1}|s_{t},a_{t},\mathbf{c}_{e})$, which depends on the embodiment, as different embodiments interact with the environment in distinct ways (due to variations in collider size and center of rotation). Fig. 2 shows example trajectories from two different embodiments starting at the same location and following the same sequence of actions. They have distinct visual observations and follow different transition dynamics—one agent moves under the table, while the other collides with it. Except where otherwise specified, we assume that all embodiments share the same discrete action space {MoveBase(±20cm), RotateBase(±$6^{\circ}$, ±$30^{\circ}$), Done}, and use robot-specific low-level controllers to execute these actions during deployment.
Figure 2: Different embodiments exhibit different behaviors. For each embodiment in these sample trajectories, the left column shows the first-person view from the main camera and the second one a third-person view of the agent –white boxes indicate the robot colliders. Embodiment A (shown on the left) has a bigger body size compared to Embodiment B (shown on the right). As a result, B can go under the table to get to the chair but A collides with the table and has to go around.
@@ 3.2 Embodiment randomization at scale
We model the body of the agent as an invisible collider box in the AI2-THOR [29] simulator. Each agent can have 1 or 2 RGB cameras placed at a random pose within the collider box. Parameters corresponding to both the body and the cameras are sampled randomly from the ranges specified in Table 1. We also modify the process of generating expert trajectories to account for the diversity of embodiments, for details see Appendix 7. Below, we detail the parameters varied in our embodiment randomization.
Collider size $(\alpha_{x},\alpha_{y},\alpha_{z})$. The agent’s body is modeled as a collider box. We use three scale factors $(\alpha_{x},\alpha_{y},\alpha_{z})$ to scale the box along $x,y,z$ axis. We sample $\alpha_{x}$ and $\alpha_{y}$ uniformly from the range $[0.2,0.5]$m and sample $\alpha_{z}$ (the height of the agent) from the range $[0.3,1.5]$m. These ranges adequately capture the variability among most robots.
Rotation center ($o_{x},o_{y},o_{z}$). These coordinates define the agent’s pivot point. While this center is typically near (0,0), it can vary across different robots. We sample $o_{x}$ from the range $[-\frac{\alpha_{x}}{3},\frac{\alpha_{x}}{3}]$ and $o_{y}$ from the range $[-\frac{\alpha_{y}}{3},\frac{\alpha_{y}}{3}]$, with the sampling ranges determined by the collider size.
Camera parameters. Each agent is equipped with two RGB cameras placed within the collider box. We randomize several camera parameters, including position, rotation, FoV, and aspect ratio. The sampling ranges for these parameters are shown in Table 1. While the first camera always faces forward, the second camera can rotate up to $360^{\circ}$ in $z$-axis, enabling it to face forward, to the sides, or backward.
For visualization purposes, we define an embodiment configuration vector $\mathbf{c}_{e}\in\mathbb{R}^{19}$ for each embodiment, representing the camera and body parameters. Fig. 1-A shows a t-SNE visualization of the vectors $\mathbf{c}_{e}$ for 30k of our random embodiments, along with the corresponding vectors for Stretch RE-1, LoCoBot, and Unitree A1. The figure also includes the egocentric view from the first camera for 10 random embodiments, and the three robots. This demonstrates that our randomization spans a wide range of possible embodiments, covering the real-world robot platforms of interest. In total, we gather 1M trajectories across 50k houses, each with a randomly sampled embodiment (more details in Appendix 10).

### v2

@@ Abstract
Modern robots vary significantly in shape, size, and sensor configurations used to perceive and interact with their environments. However, most navigation policies are embodiment-specific—a policy trained on one robot typically fails to generalize to another, even with minor changes in body size or camera viewpoint. As custom hardware becomes increasingly common, there is a growing need for a single policy that generalizes across embodiments, eliminating the need to (re-)train for each specific robot. In this paper, we introduce Ring (Robotic Indoor Navigation Generalist), an embodiment-agnostic policy that turns any mobile robot into an effective indoor semantic navigator. Trained entirely in simulation, Ring leverages large-scale randomization over robot embodiments to enable robust generalization to many real-world platforms. To support this, we augment the AI2-THOR simulator to instantiate robots with controllable configurations, varying in body size, rotation pivot point, and camera parameters. On the visual object-goal navigation task, Ring achieves strong cross-embodiment (XE) generalization—72.1% average success rate across $5$ simulated embodiments (a 16.7% absolute improvement on the Chores-$\mathbb{S}$ benchmark) and 78.9% across $4$ real-world platforms, including Stretch RE-1, LoCoBot, and Unitree Go1—matching or even surpassing embodiment-specific policies. We further deploy Ring on the RB-Y1 wheeled humanoid in a real-world kitchen environment, showcasing its out-of-the-box potential for mobile manipulation platforms.
Figure 1: We show that training on one million randomly generated embodiments in simulation (varying camera configurations, body size, and rotation pivot point) results in Ring, a generalist navigation policy that works across various robot embodiments in the real world. (A) A t-SNE visualization of embodiment parameters for 30k random agents and three real robots (we do not train on any real robot embodiment parameters). Egocentric views from the first camera are shown for 10 sample agents. (B) Ring transfers zero-shot to a wide range of embodiments in the real-world including Stretch RE-1, LoCoBot, Unitree Go1, RB-Y1 wheeled humanoid. (C) The policy displays embodiment-adaptive behavior, adjusting its navigation strategy based on its embodiment.
@@ 1 Introduction
Robot embodiments are diverse and are constantly evolving to better suit new environments and tasks. This range in body configurations—differences in size, shape, wheeled or legged locomotion, and sensor configurations—not only shapes how robots perceive the world but also how they act in it. A robot with a wide field of view (FoV) or multiple cameras can scan its surroundings quickly, while one with a narrower view might need to more actively explore a room. A small robot can squeeze through tight spaces, a low-profile one can duck under furniture, and a larger robot may need to follow more conservative routes. The influence of embodiment on behavior means a policy trained on one design, or even several, often does not perform well out of domain.
@@
There has been progress towards scalable cross-embodiment training [1, 2, 3, 4, 5]. While these methods demonstrate some transfer to unseen embodiments, they still suffer from performance degradation with relatively small changes in embodiment (e.g., camera pose modification on the same robot) [6, 7]. Potentially, this is due to these methods relying on the small amount of real-world data available in public datasets-around 20 embodiments in total [1]. Similarly, general-purpose navigation policies [8, 9, 10] are trained on datasets with relatively few embodiments (e.g., 8 robots in [9]), limiting their generalization. A more comprehensive solution is needed—one that can robustly handle the full spectrum of possible embodiments without retraining or additional adaptation.
@@
We introduce Ring, a Robotic Indoor Navigation Generalist. Ring is trained exclusively in simulation, without any use of real-world robot embodiments. In other words, all robot platforms we evaluate on (i.e., Stretch RE-1, LoCoBot, Unitree’s Go1, RB-Y1) are unseen by Ring during training. We leverage simulation to randomly sample 1 Million agent body configurations, varying the robot’s camera parameters, collider sizes, and center of rotation. Concretely, each embodiment consists of a collider box of varying dimensions and cameras with randomized parameters, placed randomly within the collider box. Fig.1-A presents a t-SNE[11] visualization of body parameters for 30k such agents. Our approach builds on the success of prior works that achieve strong real-world performance through large-scale simulation-only training [12, 13, 14]. Simulation enables training across a vast distribution of environments (150k ProcTHOR houses [15]) and objects (40k+ 3D objects in Objaverse [16]) in the AI2-THOR simulator. Extensive domain randomization on visual observations and the use of pre-trained visual encoders then allows simulation-trained policies to bridge the sim-to-real gap. We follow the training procedure in FLaRe [14], first training our policy on expert trajectories collected from $1$M randomized embodiments and subsequently fine-tuning it with on-policy reinforcement learning (RL) in the simulator.
@@
Our results demonstrate generalization to truly unseen embodiments. Ring transfers to diverse real-world embodiments without any adaptation, despite being trained entirely in simulation without access to the real robot configurations. We evaluate in a zero-shot setting across Stretch RE-1, LoCoBot, Unitree’s Go1, RB-Y1 wheeled humanoid, and even “Navigation Assistants,” where a human user captures ego-centric observations on their phone and prompts Ring to predict actions. Ring achieves $72.1\%$ average success rate in simulation (16.7% absolute improvement on Chores-$\mathbb{S}$ benchmark) and $78.9\%$ on real robot platforms—matching or even surpassing embodiment-specific policies. Ring can be further adapted to an embodiment-specialized policy with even better performance (up to 10% absolute improvement) with minimal finetuning. Ring is easy to install, and is ready for use by the community. We will release our pretrained models, code, and data.
@@ 2 Related work
Cross-embodiment. Cross-embodiment training has received substantial attention from the research community. Arguably the most representative of a large body of recent work [4, 3, 17, 5, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], Open-X-Embodiment (OXE) [30] is the fruit of a large collaboration to cover many robotic tasks, with special emphasis in manipulation. Its usage in RT-X results in a notable performance gain in emergent skill evaluations in comparison to RT-2 [31]. Despite the 1.5 million trajectories across 22 embodiments present in their dataset, the enormous cost of data collection in the real world makes further scaling challenging. CrossFormer [3] trains a transformer-based policy on 900k trajectories spanning 30 robots, drawing from OXE, navigation data from GNM [8], manipulation data from DROID [32], and additional sources. However, the limited diversity of embodiments and focus on low-level control highlight the need for denser embodiment coverage. GET-zero [33], focused on dexterous manipulation, incorporates embodiment structure via a connectivity graph to guide attention. In contrast, we generate an arbitrarily large set of randomized embodiments during training, allowing our policy to generalize zero-shot to novel embodiments without requiring access to their structure.
@@
Foundational navigation policies Following the success in recent developments for point-goal navigation [34], locomotion [35, 36, 37, 38], agile control [39], exploration [40, 41, 42], and social navigation [43], comparable results in more nuanced tasks like semantic or object-goal navigation (ObjectNav) [44, 45, 46, 47, 48, 49, 50, 51] remain elusive due to a lack of efficient exploration and semantic understanding capabilities. Recently, with powerful pretrained vision models [52, 53] and large-scale procedurally generated virtual environments [16], notable progress in end-to-end ObjectNav policy learning for specific embodiments has been achieved by means of imitation learning (IL) from shortest-path trajectories [12], RL [13], or combinations thereof [14]. In image-goal navigation, NoMaD [10], which extends ViNT [9], uses a diffusion policy to control a single embodiment. With the same goal in mind, GNM [8] trains navigation policies across 6 embodiments using IL. In contrast, our policy benefits from RL fine-tuning, improving robustness to compounding errors. Leveraging large-scale training with randomized embodiments in simulation, Ring learns a single policy that generalizes to any embodiment, including truly unseen robot platforms in the real world. Unlike NoMaD, ViNT, GNM, and Mobility VLA [54], which rely on topological map or graph reconstruction for high-level planning, our approach is fully end-to-end and can explore dynamic novel scenes without requiring an explicit map. While prior work [55, 56, 57] has explored embodiment-agnostic policies using LLMs or VLMs, these methods are limited to short-horizon navigation and single-step prediction. In contrast, Ring incorporates temporal context via a transformer decoder.
@@ 3 Ring
With the growing diversity of robots used in research labs and real-world applications, there remains a need for a policy that can operate a wide range of embodiments and transfer, in a zero- or few-shot manner, to unseen robots. We introduce Ring, a generalist policy for indoor visual navigation that learns from a broad spectrum of embodiments, trained exclusively in simulation, without any direct use of actual robot embodiments. We show that training on an extensive range of ${\sim}1$M random embodiments results in a robust navigation policy, enabling zero-shot transfer to unseen real-world robots. To train Ring, we define the space of random embodiments (Sec. 3.2), enable generation of expert trajectories for random embodiments in simulation (see Appendix 8), and use state-of-the-art architecture designs (Sec. 3.3) to train with a combination of IL and RL methods (Sec. 3.4).
3.1 Problem formulation
We define the space of possible embodiments as $E$, where each embodiment $e\in E$ is characterized by a configuration vector $\mathbf{c}_{e}$, including parameters such as camera settings, agent collider size, and center of rotation. Each task can be modeled as a Partially Observable Markov Decision Process (POMDP), denoted as $(S,A,E,O_{e},T_{e},R,L,P(s_{0}),\gamma)$, where $S$ and $A$ are the state and action spaces. The observation space $O_{e}$ varies across embodiments due to differences in camera parameters. The observation at time $t$ for embodiment $e$, $o_{t}^{e}=O_{e}(s_{t},\mathbf{c}_{e})$, is a function of both the state $s_{t}$ and embodiment parameters $\mathbf{c}_{e}$. Given an action $a_{t}$, the next state follows the transition dynamics $s_{t+1}\sim T_{e}(s_{t+1}|s_{t},a_{t},\mathbf{c}_{e})$, which depends on the embodiment (due to variations in collider size and rotation center). Fig. 2 shows trajectories from two different embodiments starting at the same location and following the same sequence of actions. They have distinct visual observations and follow different transition dynamics—one agent moves under the table, while the other collides with it.
Except where otherwise specified, we assume that all embodiments share the same discrete action space {MoveBase(±20cm), RotateBase(±$6^{\circ}$, ±$30^{\circ}$), Done}. These actions are executed using robot-specific low-level controllers during deployment. This simple and platform-agnostic action space enables effective cross-embodiment (XE) transfer, as demonstrated across both holonomic and differential-drive robots in our experiments. Investigating more expressive XE action spaces beyond this sufficient version is left for future work.
@@ 3.1 Problem formulation
Table 1: Random Embodiment Parameters. 1M embodiments are sampled from these ranges.
@@ 3.2 Embodiment randomization at scale
Figure 2: Different embodiments exhibit different behaviors. We show egocentric view from the main camera and third-person view of the 2 agents–white boxes indicate the robot colliders. Embodiment B can go under the table to get to the chair but Embodiment A collides with the table and has to go around.
We model the body of the agent as an invisible collider box in the AI2-THOR [59] simulator. Each agent can have 1 or 2 RGB cameras placed at a random pose within the collider box. Parameters corresponding to both the body and the cameras are sampled randomly from the ranges specified in Tab. 1. We also modify the process of generating expert trajectories to account for the diversity of embodiments, for details see Appendix 8. Below, we detail the parameters varied in our embodiment randomization: Collider size $(\alpha_{x},\alpha_{y},\alpha_{z})$. The agent’s body is modeled as a collider box. We use three scale factors $(\alpha_{x},\alpha_{y},\alpha_{z})$ to scale the box along $x,y,z$ axis. Rotation center ($o_{x},o_{y},o_{z}$). These coordinates define the agent’s pivot point. While this center is typically near (0,0), it can vary across different robots. We sample $o_{x}$ from the range $[-\frac{\alpha_{x}}{3},\frac{\alpha_{x}}{3}]$ and $o_{y}$ from the range $[-\frac{\alpha_{y}}{3},\frac{\alpha_{y}}{3}]$, with the sampling ranges determined by the collider size. Camera parameters. Each agent is equipped with two RGB cameras placed within the collider box. We randomize several camera parameters, including position, rotation, FoV, and aspect ratio. While the first camera always faces forward, the second camera can rotate up to $360^{\circ}$ in $z$-axis, enabling it to face forward, to the sides, or backward.
Figure 3: Ring model architecture. It accepts visual observations and a language instruction as inputs and predicts an action to execute. At RL finetuning, Ring also predicts a value estimate.